OpenStudy (anonymous):

Find the smallest positive number θ such that tanθ = -6

OpenStudy (anonymous):

i dont understand how to start this what so ever

OpenStudy (anonymous):

@amistre64 @ganeshie8 hello there i was wondering if you can help me understand this concept

ganeshie8 (ganeshie8):

you need to use ur calculator and find the angle whose tangent is -6

ganeshie8 (ganeshie8):

\[\tan (\theta) = -6\] \[\theta = \tan^{-1}(-6)\]

ganeshie8 (ganeshie8):

do you have inverse tan button in ur calculator ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got -1.4056

ganeshie8 (ganeshie8):

type in -6 and push that button

ganeshie8 (ganeshie8):

you're stuck because thats not positive

OpenStudy (anonymous):

yeahhh this wouldnt be the positive number theta ..... right its confusing

ganeshie8 (ganeshie8):

here is the hint : the period of tan is \(\pi\), so adding \(\pi\) to the angle will not change the tangent value

ganeshie8 (ganeshie8):

btw, \(\pi\) is just a number : 3.142

OpenStudy (anonymous):

mmmmhhhhh ok

ganeshie8 (ganeshie8):

simply add \(\pi\) to ur angle to make it positive

OpenStudy (anonymous):

mmmhhhh really ok ... interesting ... thank you n_n

ganeshie8 (ganeshie8):

np :) the least positive angle satisfying given equaiton would be : \[-1.4056 + 3.142 = 1.736\]